OpenStudy (anonymous):

A rectangle has a length of the fifth root of 16 inches and a width of 2 to the 1 over 5 power inches. Find the area of the rectangle.

OpenStudy (anonymous):

\[\sqrt[5]{16} \sqrt[5]{2}=\sqrt[5]{32}=2 \]